OpenStudy (anonymous):

Assume that the probability of a boy being born is the same as the probability of a girl being born. Find the probability that a family with four children will have the given composition. (Enter your answer to three decimal places.) Three boys and one girl

OpenStudy (anonymous):

\[\dbinom{4}{3}(\frac{1}{2})^4\]

OpenStudy (anonymous):

\[\dbinom{4}{3}=4\] is the number of ways you can choose 3 item out of a set of 4, so there are four differnt way you can have one girl and 4 boys g b b b b g b b b b g b b b b g

OpenStudy (anonymous):

what's with (1/2)^4?

OpenStudy (anonymous):

(1/2)^4 =1^4/2^4=1/16 2^4 =16

OpenStudy (anonymous):

there are two possibilites for each child: b or g each has probabiltty 1/2 so there are \[(\frac{1}{2})^4=\frac{1}{16}\] possible sequences of which 4 has three boy and one girl

OpenStudy (anonymous):

i understand where u get 1/16 being that there is a half chance for each and theres 4 total in question.. i entered this answer but its not right :(

OpenStudy (anonymous):

so its 4ncr3 / 2^4 = .250

OpenStudy (anonymous):

thanks!!